OpenStudy (anonymous):

someon ehelp me understand this please! The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46 centimeters. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences

OpenStudy (anonymous):

If the lengths of the sides of the triangle are given as x, x, and y. And the perimeter is given as 46, wouldn't that mean that the sum of the lengths of the sides is equal to 46? Try writing this as an equation. For the second sentence, just relate y and x in an equation. Would it be y = 4 + x or y + 4 = x ?

OpenStudy (anonymous):

y+4=x?

OpenStudy (anonymous):

If y is 4 cm longer than x then that means it requires 4 cm in addition to "x" to get y. y = x + 4. Imagine if I am 4 cm taller than you. Call me Y and you are X. Then Y = X + 4 cm!

OpenStudy (anonymous):

ooooh ok

OpenStudy (anonymous):

wait but i need two equations

OpenStudy (anonymous):

The first equation would be: x + x + y = 46. Or simply 2x + y = 46. Do you see why?

OpenStudy (anonymous):

bcs all the sides add up to that?

OpenStudy (anonymous):

Yep, and that's exactly how 'perimeter' is defined: as the sum of the sides.

OpenStudy (anonymous):

so whats the second?

OpenStudy (anonymous):

The second was : y = x + 4

OpenStudy (anonymous):

so then how would i solve the equation?

OpenStudy (anonymous):

Ah, so now you have two equations: 2x + y = 46 y = x + 4 Try the substitution method again. Let y = x + 4 and substitute that value into 2x + y = 46. This will give 2x + (x + 4) = 46. You can solve for 'x' here. Once you have x, you can solve for y.

OpenStudy (anonymous):

ok thank you once agai:=D